Write the line in slope intercept from that goes through the point (3, -2) and is parallel to y=4x+8.

Answer by math_tutor2020(3817)


Rule: parallel lines have equal slopes, but different y intercepts.

The given line
y = 4x+8
has slope m = 4.
Refer to y = mx+b form

We'll plug that m value along with the coordinates of the point (3,-2) to solve for b
y = mx+b
y = 4x+b
-2 = 4*3+b
-2 = 12+b
-2-12 = b
-14 = b
b = -14

Answer: y = 4x-14